OpenStudy (unklerhaukus):

Find \(\dfrac{\partial z}{\partial x},\dfrac{\partial z}{\partial y}\). (you may need tot use implicit differentiation) \[\sin z^2+e^u =0\qquad\qquad\text{where }u =\tan^{-1}(x-y)\]

OpenStudy (unklerhaukus):

\[u_x =\frac1{1+(x-y)^2}\\u_y =-\frac1{1+(x-y)^2}\]

OpenStudy (primeralph):

Why? Like, for real, why?

OpenStudy (unklerhaukus):

why not?

OpenStudy (primeralph):

It's one of those things you'll get eventually so I just want to know if it's a test or you really do need a solution. Because right now, I just want to watch anime.

OpenStudy (unklerhaukus):

well, i have the answers, \[z_x=-\frac{e^u}{2z\cos z^2(1+(x-y)^2)}\\ z_y=\frac{e^u}{2z\cos z^2(1+(x-y)^2)}\] but i cant work out the proper method

OpenStudy (anonymous):

well observe:$$\frac{\partial}{\partial x}[\sin z^2+e^u]=0\\2z\cos(z^2)\frac{\partial z}{\partial x}+e^u\frac{\partial u}{\partial x}=0\\2z\cos(z^2)\frac{\partial z}{\partial x}=-e^u\frac{\partial u}{\partial x}\\\frac{\partial z}{\partial x}=-\frac{e^u}{2z\cos z^2}\frac{\partial u}{\partial x}=-\frac{-e^u}{2z\cos z(1+(x-y)^2)}$$

OpenStudy (anonymous):

$$\frac{\partial}{\partial y}[\sin z^2+e^u]=0\\2z\cos(z^2)\frac{\partial z}{\partial y}+e^u\frac{\partial u}{\partial y}=0\\\frac{\partial z}{\partial y}=-\frac{e^u}{2z\cos z^2}\frac{\partial u}{\partial y}=\frac{e^u}{2z\cos z^2(1-(x-y)^2)}$$

OpenStudy (anonymous):

then just substitute in \(u=\arctan(x-y)\)... anyways this is basically the same procedure as when we do implicit ordinary differentiation

OpenStudy (unklerhaukus):

THANKYOU @oldrin.bataku